Crossett Trucking Company claims that the mean weight of its delivery trucks when they are fully loaded is 5,300 pounds and the standard deviation is 170 pounds. Assume that the population follows the normal distribution. Forty trucks are randomly selected and weighed.

Within what limits will 99 percent of the sample means occur? (Round your z-value to 2 decimal places and final answers to 1 decimal place.)

Sample means _______ to _______
Answer #1

Mean= 5300

Standard deviation= 170

N= 40

Z critical value=|z| = 2.58

Margin of error= z* sd/√N=( 2.58)(170/√40)= 69.1

5300± 69.1

Sample means 5230.9 to 5369.1